### v1

@@ Abstract
Superfund sites are locations in the United States with high levels of environmental toxicants, often resulting from industrial activity or improper waste management. Given mounting evidence linking prenatal environmental exposures to adverse birth outcomes, estimating the impact of Superfund remediation is of substantial policy relevance. A widespread approach is to fit a spatial regression, i.e., a linear regression of the outcome (e.g., birth weight) on binary treatment (e.g., indicator for Superfund site remediation) and covariates, along with a spatially structured error term to account for unmeasured spatial confounding. Despite this common practice, it remains unclear to what extent spatial regression models account for unmeasured spatial confounding in finite samples and whether such adjustments can be reformulated within a design-based framework for causal inference. To fill this knowledge gap, we introduce a weighting framework that encompasses three canonical types of spatial regression models: random effects, conditional autoregressive, and Gaussian process models. This framework yields new insights into how spatial regression models build causal contrasts between treated and control units. Specifically, we show that: 1) the spatially autocorrelated error term produces approximate balance on a hidden set of covariates, thereby adjusting for a specific class of unmeasured confounders; and 2) the error covariance structure can be equivalently expressed as regressors in a linear model. We also introduce a new average treatment effect estimator that simultaneously accounts for multiple forms of unmeasured spatial confounding, as well as diagnostics that enhance interpretability. In a study of Superfund remediation, our approach illuminates the role of design-based adjustment for confounding and provides guidance for evaluating environmental interventions in spatial settings.
Keywords: Causal Inference; Spatial Regression; Superfund Sites; Unmeasured Confounding; Weighting Methods
@@ 1.1 Superfund remediation and birth outcomes
The United States contains thousands of hazardous waste sites that potentially expose millions of Americans to toxic pollutants. More than 78 million Americans — 24% of the US population, including 17 million children and 11 million elderly individuals — live within three miles of a hazardous waste site (E.P.A. 2022). Many of these sites are legacies of historical industrial activities, including landfills, manufacturing facilities, mining sites, and processing plants. In 1980, Congress enacted the Comprehensive Environmental Response, Compensation, and Liability Act (CERCLA), granting the Environmental Protection Agency (EPA) authority to identify and remediate these sites, which became known as Superfund sites. Although cleanup efforts have been associated with reductions in congenital anomalies in infants (Currie et al. 2011) and blood lead levels in children (Klemick et al. 2020), remediation remains contentious due to high costs — often tens of millions per site — and disputes over whether the benefits justify the expense (Hamilton and Viscusi 1999).
Our goal is to estimate a causal effect of Superfund site remediation on five birth outcomes while adjusting for a rich set of sociodemographic covariates. A key challenge is unmeasured spatial confounding. Unmeasured spatial confounders are unobserved spatially structured variables that influence both treatment assignment and outcomes, potentially leading to biased effect estimates and invalid confidence intervals (Gilbert et al. 2021; Reich et al. 2021). Examples of unmeasured spatial confounders are community participation and local funding availability, which can shape both cleanup efforts and birth outcomes (Petrie 2006).
@@ 1.2 Spatial regression and spatial confounding
To adjust for potential unmeasured spatial confounding, researchers frequently introduce a mean zero spatially autocorrelated error term into a linear regression model and interpret the treatment coefficient as a causal effect, a practice that remains widespread in many disciplines. Throughout this paper, we refer to such models as “spatial regression models” (Anselin 2002). For example, a standard approach is to regress the outcome variable (e.g. birth weight) on a binary treatment variable (e.g. indicator for Superfund site remediation) and measured confounders while incorporating random intercepts at the county or state level. Alternatively, researchers use Gaussian process regression, modeling the correlation between the error terms of two Superfund sites as a function of their geographic proximity, or employ a conditional autoregressive model, which models spatial autocorrelation based on an adjacency or weight structure (Reich et al. 2021).
However, the literature offers conflicting perspectives on whether spatial regression models should be used to adjust for unmeasured spatial confounding. The prevailing view is that spatial regression should not be used to adjust for unmeasured spatial confounding, due to the bias of the treatment coefficient estimate in finite samples (Hanks et al. 2015; Hodges and Reich 2010; Khan and Berrett 2023; Narcisi et al. 2024; Nobre et al. 2021; Paciorek 2010; Page et al. 2017; Schnell and Papadogeorgou 2020). This view also arises from the recognition that a spatial regression model is inherently misspecified when there is unmeasured spatial confounding. Specifically, spatially structured error is typically modeled independently of treatment, under the assumption that it does not capture unmeasured spatial confounding — which, if present, would induce dependence between the error term and the treatment (Hodges and Reich 2010).
Recently, Gilbert et al. (2025) showed that, under certain conditions, spatial regression models with Gaussian process covariance yield consistent treatment effect estimates under unmeasured confounding that is continuous in space. Although the bias of treatment coefficient is non-zero in finite samples, the authors show that it vanishes asymptotically. While this result is of substantial interest, it remains uncertain whether similar consistency guarantees extend to other commonly used spatial regression models — such as random effects models and conditional autoregressive models — or what implications this has for the magnitude of unmeasured spatial confounding bias in finite samples.
@@ 1.3 Contribution and outline
The aforementioned literature is subject to limitations. Each approach studies the bias of spatial regression estimators under the restrictive assumption of an outcome model with a linear effect of treatment and without any interactions. Additionally, Gilbert et al. (2025); Nobre et al. (2021); Paciorek (2010); Page et al. (2017) specify a particular form for the error covariance matrix, such as a Gaussian process covariance matrix. The impact of this choice on unmeasured confounding bias is not explored. Dupont et al. (2023); Khan and Berrett (2023); Narcisi et al. (2024) leave the error covariance matrix in a general form, but their bias results — expressed qualitatively in terms of the relationship between the unmeasured confounder and the eigenstructure of the precision matrix — are not straightforward to interpret. Finally, Narcisi et al. (2024); Nobre et al. (2021); Paciorek (2010); Page et al. (2017) assess the bias while imposing strong distributional assumptions on the unmeasured confounder $U$, such as multivariate normality.
@@
In general, the adjustment mechanism by which a spatial regression model accounts for unmeasured spatial confounding remains to be understood, particularly in finite samples, leaving open questions about the hypothetical experiment and the causal contrast it defines in space. In this paper, we provide an exact finite-sample analysis of spatial regression models in binary treatment settings. Our contributions are fourfold. First, we introduce a general weighting framework for causal inference that encompasses three canonical types of spatial regression models: random effects models, conditional autoregressive models, and Gaussian process models. We show that generalized least squares (GLS) estimators derived from spatial regression models are equivalent to minimal dispersion balancing weights estimators, which minimize a measure of weight dispersion subject to covariate balance requirements (Bruns-Smith et al. 2023; Chattopadhyay and Zubizarreta 2023; Kline 2011; Robins et al. 2007). Our equivalence results are the first to show that adjustments made by GLS regression approximately balance the means of a hidden set of additional covariates. The hidden set of additional covariates corresponds to the eigenvectors of the error covariance matrix. This perspective is related to the work of Hodges and Reich (2010), who showed that the spatially correlated error term in a linear model can be represented through additional canonical regressors with random coefficients. Importantly, we further demonstrate that the resulting GLS regression is equivalent to a form of penalized regression, and we provide a mathematical programming representation that characterizes the exact nature of this regression adjustment.
@@
Second, we characterize the class of unmeasured confounders that the three spatial regression models are designed to adjust for. We do so by relating the unmeasured confounding bias due to $U$ to its Moran’s I statistic, which is a measure of spatial autocorrelation of $U$. This mathematical characterization is accompanied by three visual diagnostics to complement traditional weighting diagnostics: a graph of the maximum unmeasured confounding bias in finite sample as a function of the unmeasured confounder’s Moran’s I, illustrative examples of unmeasured confounders that are approximately adjusted for, and a graph of implied weights.
@@
Third, we propose a more general spatial weighting estimator under a causal inference framework that overcomes many of the limitations of spatial regression. Our proposed estimator accounts for multiple forms of unmeasured confounders with spatial autocorrelation, while accommodating nonlinearities and treatment effect heterogeneity by the unmeasured confounder in the outcome model. We derive bounds on the finite-sample bias of our estimator and evaluate its performance in simulation.
Fourth, we apply our proposed spatial weighting estimator to study the impact of Superfund site remediation on five birth outcomes. Among other insights, our approach reveals how the causal contrast of remediated and non-remediated Superfund sites is built across the geography of the United States.
@@
This paper proceeds as follows. Section 2 describes the dataset on Superfund sites and birth outcomes. Section 3 introduces three common examples of spatial regression models, including models with random intercepts at the state level, conditional autoregressive models, and Gaussian process models. Section 4 establishes a general weighting framework that includes GLS estimators derived from spatial regression models as a special case. Section 5 relates the unmeasured confounding bias due to $U$ to its Moran’s I statistic and develops new diagnostics illustrating this result. Section 6 proposes an alternative estimator to spatial regression and derives its finite-sample properties in the presence of unmeasured spatial confounding. Section 7 evaluates the performance of our estimator in simulation and Section 8 uses it to estimate a causal effect of Superfund site remediation on five birth outcomes in New Jersey.
2 Superfund site remediation in the United States
In this study, we examine the impact of Superfund site remediation on five birth outcomes among the sites that were cleaned. The dataset is publicly available and fully reproducible code is accessible at https://github.com/NSAPH-Projects/spatial_regression_weighting.
The dataset includes $n=1583$ Superfund sites in the United States that were listed on the National Priorities List (NPL) as of 2001. The binary treatment variable $Z_{i}\in\{0,1\}$ indicates whether the $i$th Superfund site was remediated and removed from the NPL between 2001 and 2015. Figure 1 plots binary treatment.
@@ 2 Superfund site remediation in the United States
Figure 1: Binary Treatment. There were $n=1583$ Superfund sites on the National Priority List on January 1, 2001. During the fifteen-year period between January 1, 2001 and December 31, 2015, $n_{1}=159$ sites were cleaned up and removed from the National Priority List.
@@
We consider ten sociodemographic confounders derived from the 2000 Decennial Census, measured within 2 kilometer buffers around each site. These covariates, originally obtained at the census tract-level and subsequently summarized to the buffer-level, include population density, percentage of individuals identifying as Hispanic, Black, Indigenous, or Asian, median household income, median house value, percentage below the poverty line, percentage with a high school diploma, and the median year of housing construction. We additionally include site score as a covariate — a numerical rating assigned to each site by the EPA during a hazardous risk assessment. Let $\bm{X}_{i}\in\mathbb{R}^{12}$ denote the vector of observed covariates for site $i$, including an intercept.
We consider five birth outcomes measured within $2$ kilometers of site $i$ during the period 2016 to 2020. Importantly, this paper focuses on the design stage of spatial regression and outcome data is not used until Section 8.
Under the Stable Unit Treatment Value Assumption (SUTVA), let $\{Y_{i}(1),Y_{i}(0)\}$ be the potential outcomes under treatment and control, respectively, such that $Y_{i}=Z_{i}Y_{i}(1)+(1-Z_{i})Y_{i}(0)$. We assume that the coefficient estimate of treatment obtained from a spatial regression model is interpreted by the researcher, implicitly or explicitly, as the average treatment effect on the treated (ATT). This estimand is defined as $\tau_{ATT}=\mathbb{E}(Y_{i}(1)-Y_{i}(0)|Z_{i}=1).$
@@ 3 Three common spatial regression models
In this section, we describe a general form of a spatial regression model and introduce three widely used classes of such models for estimating causal effects in spatial settings. For clarity of notation, let $\bm{X}=(\bm{X}_{1},\ldots,\bm{X}_{n})^{\top}\in\mathbb{R}^{n\times 12}$ denote the matrix of observed covariates, $\bm{Z}=(Z_{1},\ldots,Z_{n})^{\top}\in\mathbb{R}^{n}$ represent the vector of treatment assignments, and $\bm{Y}=(Y_{1},\ldots,Y_{n})^{\top}\in\mathbb{R}^{n}$ be the vector of outcomes.

### v2

@@ Abstract
A key challenge in environmental health research is unmeasured spatial confounding, driven by unobserved spatially structured variables that influence both treatment and outcome. A common approach is to fit a spatial regression that models the outcome as a linear function of treatment and covariates, with a spatially structured error term to account for unmeasured spatial confounding. However, it remains unclear to what extent spatial regression actually accounts for such forms of confounding in finite samples, and whether this regression adjustment can be reformulated from a design-based perspective. Motivated by an observational study on the effect of Superfund site remediation on birth outcomes, we present a weighting framework for causal inference that unifies three canonical classes of spatial regression models—random effects, conditional autoregressive, and Gaussian process models—and reveals how they implicitly construct causal contrasts across space. Specifically, we show that: (i) the spatial error term induces approximate balance on a latent set of covariates and therefore adjusts for a specific form of unmeasured confounding; and (ii) the covariance structure of the spatial error can be equivalently represented as regressors in a linear model. Building on these insights, we introduce a new estimator that jointly addresses multiple forms of unmeasured spatial confounding and develop visual diagnostics. Using our new estimator, we find evidence of a small but beneficial effect of remediation on the percentage of small vulnerable newborns.
Keywords: Causal Inference; Spatial Regression; Superfund Site; Unmeasured Confounding; Weighting Method
@@ 1.1 Superfund remediation and birth outcomes
The United States contains thousands of hazardous waste sites that potentially expose millions of Americans to toxic pollutants. More than 78 million Americans—24% of the US population, including 17 million children and 11 million elderly individuals—live within three miles of a hazardous waste site (E.P.A. 2022) . Many of these sites are legacies of historical industrial activities, including landfills, manufacturing facilities, mining sites, and processing plants. In 1980, Congress enacted the Comprehensive Environmental Response, Compensation, and Liability Act (CERCLA), granting the Environmental Protection Agency (EPA) authority to identify and remediate these sites, which became known as Superfund sites. Although cleanup efforts have been associated with reductions in congenital anomalies in infants (Currie et al. 2011) and blood lead levels in children (Klemick et al. 2020), remediation remains contentious due to high costs—often tens of millions of dollars per site—and disputes over whether the benefits justify the expense (Hamilton and Viscusi 1999).
Our goal is to estimate the causal effect of Superfund site remediation on two birth outcomes, while adjusting for a rich set of sociodemographic covariates. A key challenge is unmeasured spatial confounding. Unmeasured spatial confounders are unobserved spatially structured variables that influence both treatment assignment and outcomes, potentially leading to biased effect estimates and invalid confidence intervals (Gilbert et al. 2021; Reich et al. 2021). Examples include community participation and local funding availability, which can shape both cleanup efforts and birth outcomes (Petrie 2006).
@@ 1.2 Spatial regression and spatial confounding
To address potential unmeasured spatial confounding, researchers frequently introduce a mean zero spatially autocorrelated error term into a linear regression model and interpret the treatment coefficient as a causal effect, a practice that remains widespread in many disciplines. Throughout this paper, we refer to such models as “spatial regression models” (Anselin 2002). For example, a standard approach is to regress the outcome variable (e.g., birth weight) on a binary treatment variable (e.g., Superfund site remediation) and measured confounders, while incorporating random intercepts at the county or state level. Alternatively, researchers use Gaussian process regression, modeling the correlation between the error terms of two Superfund sites as a function of their geographic proximity, or employ a conditional autoregressive model, which models spatial autocorrelation based on an adjacency or weight structure (Reich et al. 2021).
The literature, however, offers conflicting perspectives on whether spatial regression models should be used to adjust for unmeasured spatial confounding. A prevailing view is that spatial regression is inappropriate for this purpose, because the estimated treatment coefficient can be biased in finite samples (Hanks et al. 2015; Hodges and Reich 2010; Khan and Berrett 2023; Narcisi et al. 2024; Nobre et al. 2021; Paciorek 2010; Page et al. 2017; Schnell and Papadogeorgou 2020). This view also arises from the fact that a spatial regression model is inherently misspecified when there is unmeasured spatial confounding. Specifically, spatially structured error is typically modeled independently of treatment under the assumption that it does not capture unmeasured spatial confounding, which, if present, would induce dependence between the error term and the treatment (Hodges and Reich 2010).
Recently, Gilbert et al. (2025) showed that, under certain conditions, spatial regression models with Gaussian process covariance yield consistent treatment effect estimates under unmeasured confounding that is continuous in space. Although the bias of the treatment coefficient is non-zero in finite samples, the authors show that it vanishes asymptotically. While this result is of substantial interest, it remains uncertain whether similar consistency guarantees extend to other commonly used spatial regression models—such as random effects models and conditional autoregressive models—or what implications this has for the magnitude of unmeasured spatial confounding bias in finite samples.
@@ 1.3 Contribution and outline
Although this literature has advanced our understanding of spatial regression under unmeasured confounding, several critical aspects remain underexplored. First, most bias analyses focus primarily on linear outcome models with constant treatment effects and no treatment-covariate interactions. Second, while studies such as Gilbert et al. (2025) and Paciorek (2010) use particular forms for the error covariance matrix, such as a Gaussian process structure, it is worth exploring how these choices adjust for bias in finite samples. This question extends to other error structures, such as random effects and conditional autoregressive models. Third, although Dupont et al. (2023), Khan and Berrett (2023), and Narcisi et al. (2024) accommodate a general error covariance matrix, they characterize bias through the eigenstructure of the precision matrix rather than through more interpretable measures of spatial autocorrelation. Finally, current methods often rely on strong distributional assumptions regarding the unmeasured confounder $U$, such as multivariate normality.
@@
More broadly, the adjustment mechanism by which spatial regression accounts for unmeasured spatial confounding remains to be understood from a design-based standpoint, leaving open questions about the hypothetical experiment and the causal contrast these models implicitly define in space. In this paper, we provide an exact finite-sample analysis of spatial regression models in binary treatment settings. Our contributions are fourfold. First, we introduce a general weighting framework for causal inference with spatial data that encompasses three canonical types of spatial regression models: random effects models, conditional autoregressive models, and Gaussian process models. We establish that generalized least squares (GLS) estimators of the treatment effect derived from spatial regression models are equivalent to minimal dispersion approximately balancing weights estimators, which minimize a measure of weight dispersion subject to specific covariate balance requirements (Bruns-Smith et al. 2023; Chattopadhyay and Zubizarreta 2023; Robins et al. 2007). To our knowledge, these are the first results to show that a GLS estimator is equivalent to a weighting estimator that achieves exact balance on the means of measured covariates and approximate balance on an additional latent set of covariates—specifically, the eigenvectors of the error covariance matrix. This result elucidates how the error structure adjusts estimates to account for spatial patterns. This perspective is related to the work of Hodges and Reich (2010), who showed that the spatially correlated error term in a linear model can be represented through additional canonical regressors with random coefficients. Importantly, we further demonstrate that the resulting GLS regression is equivalent to a form of penalized regression, and we provide a mathematical programming representation that characterizes the exact nature of this regression adjustment.
@@
Second, we characterize the class of unmeasured confounders that these three spatial regression models are designed to adjust for. We do so by linking the unmeasured confounding bias to the confounder’s Moran’s I statistic, a standard measure of spatial autocorrelation. This mathematical characterization is accompanied by three visual diagnostics to complement traditional weighting diagnostics: a map of the implied weights, visualizations of unmeasured confounders that are approximately balanced, and a plot of the maximum confounding bias as a function of the unmeasured confounder’s Moran’s I.
@@
Third, we propose a more general spatial weighting estimator that circumvents key limitations of traditional spatial regression. Unlike standard approaches, our estimator accounts for multiple forms of spatially autocorrelated unmeasured confounding while accommodating nonlinear outcome models and treatment effect heterogeneity. We derive bounds on the finite-sample bias of our estimator and evaluate its performance through simulation.
Fourth, we apply our proposed spatial weighting estimator to study the impact of Superfund site remediation on birth outcomes. To our knowledge, this study represents the first national-scale linkage of Medicaid claims with EPA remediation data to evaluate health impacts of Superfund cleanup. Our approach reveals how the causal contrast of remediated and non-remediated Superfund sites is constructed across the geography of the United States, making explicit the implicit geography of the comparison. Our results suggest that cleanup efforts between 1991 and 2015 had a small but statistically significant beneficial effect on the rate of small, vulnerable newborns during 2016–2018.
@@
The remainder of this paper is organized as follows. Section 2 provides institutional and scientific background on the Superfund program and describes the linked dataset used for analysis. Section 3 then reviews three canonical types of spatial regression models: random effects, conditional autoregressive, and Gaussian process models. Section 4 establishes a general weighting framework that encompasses GLS estimators derived from these spatial regression models as a special case. Section 5 characterizes unmeasured confounding bias through Moran’s I and introduces new visual diagnostics. Building on this, Section 6 proposes an alternative estimator to spatial regression and describes its finite-sample properties in the presence of unmeasured spatial confounding. Simulation results are detailed in the Supplementary Materials. Section 7 applies our approach to estimate the causal effect of Superfund site remediation on Medicaid-covered birth outcomes. Section 8 concludes with practical guidance for evaluating environmental interventions in spatial settings.
2 Superfund remediation in the United States
2.1 Background
Remediation of Superfund sites is a federally managed process that aims to identify, prioritize, and clean hazardous waste sites to protect human health and the environment. Remediation typically proceeds through several stages, beginning with site identification and preliminary assessment, followed by placement on the National Priorities List (NPL), detailed investigation of contamination, design and implementation of remedial actions, and finally site deletion once cleanup goals are achieved.
Superfund site remediation may improve birth outcomes in nearby areas by reducing maternal exposure to contaminants—such as arsenic, lead, other metals, volatile organic compounds, and persistent organic pollutants—that have been linked to adverse pregnancy and birth outcomes, or by interrupting these exposure pathways (Ouidir et al. 2020; Perera et al. 2003; Porpora et al. 2019; Stillerman et al. 2008). However, quantitative evidence on the causal effects of remediation on maternal and infant health remains limited. Currie et al. (2011) report that remediation may reduce the incidence of congenital anomalies by 20–25% but find no significant effects on low birth weight, preterm birth, or infant mortality. Most prior studies have examined associations between residential proximity to contaminated sites and birth outcomes rather than the causal impact of remediation itself (Brender et al. 2011; Henn et al. 2016; Keeler et al. 2023; Kihal-Talantikite et al. 2017; Langlois et al. 2009).
Evaluating the causal effect of remediation is complicated by several variables that influence both site remediation and birth outcomes (confounders) that are difficult to measure or obtain. One example is the severity of the contamination itself: sites that undergo cleanup are often those with less severe or more tractable contamination, and the severity of contamination likely affects the severity of adverse outcomes. A second example is local funding availability. Remediation is typically financed by potentially responsible parties, but when such parties cannot be identified or are unable to pay, funding may come from a combination of the Superfund trust fund, taxpayers, states, or an excise tax on chemical manufacturers, making funding availability complex (Kiel and Zabel 2001). Local funding availability may also correlate with community wealth and access to healthcare. While both of these variables are challenging to measure, they are likely to be spatially autocorrelated, providing an opportunity to account for them through a spatial confounding adjustment.
2.2 Data
In this study, we evaluate the impact of Superfund site remediation on birth outcomes through a novel synthesis of publicly available EPA and high-resolution census data with Medicaid birth claims from 2016–2018, which cover about 40% of all U.S. births during this period, while accounting for potential bias from unmeasured spatial confounders. The dataset includes $n=1429$ Superfund sites in the United States currently on the National Priorities List (NPL) as of October 30, 2025, excluding those for which remedial action had started before January 1, 1991. The binary treatment variable indicates whether a Superfund site was remediated between 1991 and 2015. More specifically, we define a site as treated if the site was deleted from the NPL before December 31, 2015, and control otherwise. Figure 1 plots binary treatment and Figure 11 in the Supplementary Material illustrates the definition of binary treatment.
@@ 2.2 Data
Figure 1: As of January 1, 1991, remediation action had not started at $n=1429$ Superfund sites. During the fifteen year period between January 1, 1991 and December 31, 2015, $n_{1}=256$ sites (blue triangles) were remediated and deleted from the National Priorities List.
@@
We consider eleven sociodemographic confounders derived from the 1990 Decennial Census, measured within each site’s boundary and the surrounding 2-kilometer buffer. The buffer size is motivated by prior work examining infant health near Superfund sites (Currie et al. 2011). The census covariates, originally obtained at the census tract-level and subsequently summarized to the buffer-level, include population density, percentage of individuals identifying as Hispanic, Black, Indigenous, or Asian, percent owner occupied, median household income, median house value, percentage below the poverty line, percentage with a high school diploma, and the median year of housing construction (U.S. Census Bureau 1991; Schroeder et al. 2025). We additionally include five binary indicators: two indicating whether at least one metal or at least one volatile organic compound (VOC) was present, and three indicating whether at least one contaminant was detected in soil or other solids, water, or gas . Finally, we include site score, a numerical rating assigned to each site by the EPA during a hazardous risk assessment. Treatment and covariate data are publicly available on Harvard Dataverse at https://doi.org/10.7910/DVN/EKSCCU and fully reproducible code is accessible at https://github.com/NSAPH-Projects/spatial_regression_weighting.
We consider two birth outcomes measured within each site’s boundary and the surrounding $2$-kilometer buffer during the period 2016 to 2018. Importantly, this paper focuses on the design stage of spatial regression, and the outcome data are not used until Section 7. The two outcomes are (i) the percentage of newborns classified as small vulnerable—defined as preterm, small for gestational age, or low birth weight (Ashorn et al. 2023)—and (ii) the percentage of newborns with congenital anomalies. Both variables were obtained at the zip code tabulation area (ZCTA) level from Medicaid T-MSIS Analytic File Inpatient Claims (Centers for Medicare & Medicaid Services (2022) CMS). These ZCTA-level statistics were aggregated to the Superfund site and $2$-kilometer buffer areas (see Supplementary Section I for more details).
@@ 3 Three common spatial regression models
In this section, we describe a general form of a spatial regression model and introduce three widely used classes of such models for estimating causal effects in spatial settings. Write $Z_{i}\in\{0,1\}$ for the binary treatment variable, $\bm{X}_{i}\in\mathbb{R}^{18}$ for the vector of $17$ observed covariates and the intercept, and $Y_{i}\in\mathbb{R}$ for one of the two birth outcomes for site $i$. Further write $\bm{Z}=(Z_{1},\ldots,Z_{n})^{\top}\in\mathbb{R}^{n}$ for the vector of binary treatments, $\bm{X}=(\bm{X}_{1},\ldots,\bm{X}_{n})^{\top}\in\mathbb{R}^{n\times 18}$ for the matrix of observed covariates, and $\bm{Y}=(Y_{1},\ldots,Y_{n})^{\top}\in\mathbb{R}^{n}$ for the vector of outcomes. Under the Stable Unit Treatment Value Assumption (SUTVA), define $\{Y_{i}(1),Y_{i}(0)\}$ as the potential outcomes under treatment and control, respectively, such that $Y_{i}=Z_{i}Y_{i}(1)+(1-Z_{i})Y_{i}(0)$.